OpenStudy (anonymous):

18 people were poled and 10 said they have shopped at Wal-Mart and 8 said they have shopped at Target. 4 said they shopped at both. What is the probability that the group shopped at Wal-Mart or Target?.

OpenStudy (anonymous):

We know that 4 people shopped at both, so if we subtract 4 from the people who shopped at Walmart and Target, we get the number of people who shopped only at Walmart and Target. So only Walmart = 10 - 4 = 6 Only Target = 8 - 4 = 4 So we're left with 4 people who didn't shop at either. So the probability of someone in the group shopping at either only Walmart or Target is: # of only Target) + # of only WalMart/ total # which is: (6+4)/18 = 10/18